# Pretraining Frame Preservation in Autoregressive Video Memory Compression

### v3

@@ 4.2 Ablation Study
Table 1: Quantitative results on compression structures. We show numerical tests with different ablative compression architectures. * The “Large Patchifier” is technically equivalent to [78].
@@
Compression model architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [78] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
@@
Figure 6: Influence of memory compression model pretraining. We present results with or without the pretraining of the memory compression model. The input is the same 20-second history video, and we visualize the middle frames in the output shots.
@@ 4.5 Alternative Architecture
Figure 9: Adding cross-attention enhancing. We show that connecting the compression model and the DiT with cross-attention layers can improve consistency in difficult cases.
@@
Figure 10: Using multiple memory compression models. We show that using multiple compression models at the same time with different compression patterns, e.g., higher temporal and higher spatial, can facilitate detail consistency in difficult cases, at the cost of doubling the context length.
Combining multiple compression models. As shown in Fig. 10, another enhancement is to use multiple memory encoders simultaneously. For instance, a standard $4\times 4\times 2$ encoder can be paired with a $2\times 2\times 8$ encoder, which prioritizes temporal compression while preserving more spatial details. Using multiple encoders allows the model to capture different aspects of the history. This architecture offers advantages such as preserving fine-grained details, like text on newspapers, magazines, or billboards, at the cost of an increased context length.
@@ 5 Conclusion
In this work, we discuss the fundamental trade-off between context length and quality in autoregressive video generation as a critical challenge for creating long, coherent narratives. We introduced a framework pretraining a memory compression model with an explicit objective: high-fidelity retrieval of arbitrary frames from a long video history. This pretraining phase imposes the model to learn a compact yet detail-rich representation of the video context, effectively preserving high-frequency details. By subsequently fine-tuning this pretrained encoder within an autoregressive video diffusion system, we developed a practical and efficient model capable of handling long-term dependencies, such as maintaining character identity and scene consistency.

### v5

@@ 4.2 Ablation Study
Table 1: Quantitative results on encoder architectures. We show numerical tests with different ablative architectures. * The “Large Patchifier” is technically equivalent to [77].
@@
Encoder architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [77] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
@@
Figure 6: Influence of history encoder pretraining. We present results with or without the pretraining of the history encoder. The input is the same history video, and we visualize the middle frames in the output shots.
@@ 4.5 Alternative Architecture
Figure 9: Adding cross-attention enhancing. We show that connecting the history encoder and the DiT with cross-attention layers can improve consistency in difficult cases.
@@
Figure 10: Using multiple history encoders. We show that using multiple encoders at the same time with different compression patterns, e.g., higher temporal and higher spatial, can facilitate detail consistency in difficult cases, at the cost of doubling the context length.
Combining multiple encoders. As shown in Fig. 10, another enhancement is to use multiple history encoders simultaneously. For instance, a standard $4\times 4\times 2$ encoder can be paired with a $2\times 2\times 8$ encoder, which prioritizes temporal compression while preserving more spatial details. Using multiple encoders allows the model to capture different aspects of the history. This architecture offers advantages such as preserving fine-grained details, like text on newspapers, magazines, or billboards, at the cost of an increased context length.
@@ 5 Conclusion
In this work, we study the challenge of encoding long video histories into lightweight context for autoregressive video generation, targeting personal and local workflows with limited compute. We present a history encoder pretrained with a frame query objective: the encoder learns to attend to arbitrary frames from a long video history and encode them into a lightweight embedding. By subsequently finetuning this pretrained encoder within an autoregressive video diffusion system, the resulting model maintains long-range content consistency, including character identity, clothing, and scene layout. We evaluate the framework with ablative settings and discuss the trade-offs of different architecture designs.